OpenStudy (anonymous):

A right triangle has a hypotenuse of length 10 and a leg of length 7. What is the area of the triangle?

OpenStudy (anonymous):

a=35?

OpenStudy (ash2326):

 Find the missing side

OpenStudy (ash2326):

@marcoduuuh Are you trying?

OpenStudy (anonymous):

yes!

OpenStudy (ash2326):

Find the side then use the area formula, \[\frac 1 2 \times base\times \ height\]

OpenStudy (anonymous):

hey just an idea here take half of seven and than half of 10 and times them together

OpenStudy (anonymous):

I got 35?

OpenStudy (anonymous):

1/2 of 7 is 3.5 and half of 10 is 5 so what is the answer?

OpenStudy (ash2326):

 You need to find the height first, formula doesn't involve hypotenuse's length

OpenStudy (anonymous):

17.5!

OpenStudy (ash2326):

@marcoduuuh Can you use Pythagoras to find height?

OpenStudy (anonymous):

better use a2+b2=c2 may be that would work here

OpenStudy (anonymous):

that's Pythagoras

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hey ash am I missing something here??

OpenStudy (anonymous):

10^2+7^2= 149. sqrt of 149 is: 12.2065556

OpenStudy (anonymous):

does this problem have answer choices or anything??

OpenStudy (anonymous):

one sec checking something here

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wait I know now

OpenStudy (anonymous):

opp/ adj what is the function?

OpenStudy (anonymous):

you still there?